# Constructive agents nullify the ability of destructive agents to foster cooperation in public goods games

### v2

@@ IV Discussion
Although constructive agents themselves cannot directly alter the evolutionary outcomes of the population, together with destructive agents, they can engage in a cyclic dominance induced by destructive agents to sustain cooperation (see Figure 3). Moreover, constructive agents act as a buffer, allowing cooperation to be maintained even when the damage from destructive agents is larger. However, as the benefits provided by constructive agents increase, the efficiency of destructive agents in promoting cooperation decreases. When the benefits provided by constructive agents exceed the “harm” caused by destructive agents to individuals, it breaks the cyclic dominance and weakens the ability of destructive agents to sustain cooperation (see Figure 4). Furthermore, an increased mutation rate can weaken the peak cooperation frequency in cyclic dominance. Conversely, in the limit of rare mutation rates, the phenomena of cyclic dominance among strategies completely disappear, with the defective strategy being the unique Nash equilibrium. Neither destructive nor constructive agents derive any benefits from a common pool of funds. However, relying solely on altruistic behavior from constructive agents is insufficient to support the establishment of cooperation in one-shot games. Importantly, the altruistic behavior of constructive agents provides positive incentives for others in society, while the behavior of destructive agents generates negative incentives. By simultaneously considering these positive and negative incentive behaviors Balliet, Mulder, and Van Lange (2011); Szolnoki and Perc (2010, 2017), the dynamics of cooperation maintenance are altered. Therefore, our study highlights the importance of balancing the interaction of positive and negative incentive behaviors.
@@
In our study, constructive and destructive agents represent two typical orientations towards the interests of others: benefiting others and causing harm to others. However, individual self-interest considerations have been overlooked. Considering that self-interest is a characteristic of rational human behavior, an integration of individual actions build upon two dimensions of behavioral orientation within the social value orientation framework Bogaert, Boone, and Declerck (2008)–preference for one’s own gains and preference for opponent’s gains–will contribute to a more comprehensive understanding of cooperation emergence and promotion under diverse behavioral frameworks. On the other hand, in this challenging one-shot game scenario, cooperation cannot be sustained by relying solely on the constructive agents. In conjunction with reciprocity mechanisms, can the altruistic behaviors exhibited by constructive agents, promote the development of cooperation? Furthermore, by integrating behavioral information transmission, analyzing the biased strategies of constructive and destructive agents towards defectors or cooperators becomes feasible, aiding in a deeper understanding of how diverse behaviors affect the cooperation dynamics. Future research focusing on direct reciprocity, indirect reciprocity, and network reciprocity mechanisms will offer valuable insights into the complex interplay between diverse individual value orientations and the maintenance of cooperation.
Acknowledgements.
@@
We acknowledge the support provided by (i) the National Natural Science Foundation of China (Grant Nos.11931015, 12271471, and 12161089), Major Program of National Fund of Philosophy and Social Science of China (Grant Nos. 22&ZD158 and 22VRCO49) to L.S.; (ii) JSPS KAKENHI (Grant No. JP 23H03499) to C. S.; (iii) the Grant-in-Aid for Scientific Research from JSPS, Japan, KAKENHI (Grant Nos. JP 20H02314 and JP 23H03499) awarded to J.T.; (iv) China Scholarship Council (Grant No. 202308530309) and Yunnan Provincial Department of Education Science Research Fund Project (Grant No. 2024Y503) to Z.H.; and (v) Yunnan Provincial Department of Education Science Research Fund Project (Grant No. 2024Y502) to Y.D.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
@@
Data Availability Statement
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Appendix A Finite populations: invasion analysis among cooperation, defection and constructive agents
@@
$\begin{gathered}P_{X}=\pi_{X}(1X,\ (N-1)Y).\end{gathered}$ (10)
@@
$\begin{gathered}P_{Y}=\pi_{Y}(1X,(N-1)Y)\cdot p_{1}\\ +\pi_{Y}(0X,NY)\cdot p_{2},\end{gathered}$ (11)
@@
$\begin{split}p_{1}&=\frac{\binom{M-2}{N-2}\binom{1}{1}}{\binom{M-1}{N-1}}\\ &=\frac{N-1}{M-1},\end{split}$ (12)
$p_{2}$ $=\frac{\binom{M-2}{N-1}\binom{1}{0}}{\binom{M-1}{N-1}}$ (13)
@@
$r<N\cdot\frac{M-1}{M-N}.$ (15)
In the limit $M\to\infty$ , the above condition is reduced to $r<N$, thereby encompassing the condition for the social dilemma: $0<r<N$. In the subsequent analysis, we will proceed under the assumption that inequality (15) holds.
@@
$0<r<1-\frac{d_{2}}{M-1},$ (18)
strategy $CA$ can invade the population of strategy $C$ iff condition (18) holds.
@@
$r>1-\frac{d_{2}}{M-1},$ (19)
when $r$ satisfies condition (19), the strategy $CA$ can never invade the population of strategy $C$
@@
$r>1-(N-1)\cdot d_{2},$ (21)
strategy $C$ will invade the population of strategy $CA$ iff condition (21) holds.
@@
Appendix B Finite populations: invasion analysis of a simplified three-strategy model involving cooperation, defection and integrated exiters
@@
If $P_{D}>P_{C}$, strategy $D$ successfully invades the population of strategy $C$ , consistent with (15), i.e., when
$r<N\cdot\frac{M-1}{M-N}.$ (26)
@@
Since inequality (26) includes the condition $0<r<N$, it follows that $P_{C}<P_{D}$, indicating that strategy $C$ will never invade the population of strategy $D$.
@@
In the above equation (34), if $d_{2}-d_{1}>0$, then $P_{D}>P_{E}$, and strategy $D$ will invade the population of strategy $E$. Conversely, if $d_{2}-d_{1}<0$, then $P_{D}<P_{E}$, and strategy $D$ does not invade the population of strategy $E$. When $d_{2}=d_{1}$, strategy $D$ and strategy $E$ cannot invade each other.
@@
Similarly, in equation (35), if $d_{2}-d_{1}<0$, then $P_{E}>P_{D}$, and strategy $E$ will invade the population of strategy $D$. Conversely, if $d_{2}-d_{1}>0$, then $P_{E}<P_{D}$, and strategy $E$ does not invade the population of strategy $D$. When $d_{2}=d_{1}$, strategy $D$ and strategy $E$ cannot invade each other.
@@
However, when $d_{2}-d_{1}=0$, it follows from (34) and (35) that $PE=PD=0$, indicating that strategies $D$ and $E$ cannot invade each other, and there is no competition between them. For strategies $C$ and $E$, according to (28) and (31), we have $P_{C}=r-1$ and $P_{E}=0$. Therefore, when $r>1$, strategy $E$ does not invade the population of strategy $C$; instead, strategy $C$ invades the population of strategy $E$. Additionally, based on (25) and (27), strategy $D$ always invades the population of strategy $C$. Thus, this leads to $E\rightarrow C\rightarrow D$, resulting in the dominance of strategy $D$. Conversely, when $r<1$, strategy $C$ does not invade strategy $E$, while strategy $E$ invades strategy $C$. Combined with (25) and (27), since strategy $D$ always invades strategy $C$, both strategies $E$ and $D$ invade strategy $C$, ultimately leading to the coexistence of strategies $E$ and $D$.

### v3

@@ IV Discussion
Although constructive agents themselves cannot directly alter the evolutionary outcomes of the population, together with destructive agents, they can engage in a cyclic dominance induced by destructive agents to sustain cooperation (see Figure 3). Moreover, constructive agents act as a buffer, allowing cooperation to be maintained even when the damage from destructive agents is larger. However, as the benefits provided by constructive agents increase, the efficiency of destructive agents in promoting cooperation decreases. When the benefits provided by constructive agents exceed the “harm” caused by destructive agents to individuals, it breaks the cyclic dominance and weakens the ability of destructive agents to sustain cooperation (see Figure 4). Furthermore, an increased mutation rate can weaken the peak cooperation frequency in cyclic dominance. Conversely, in the limit of rare mutation rates, the phenomena of cyclic dominance among strategies completely disappear, with the defective strategy being the unique Nash equilibrium. Neither destructive nor constructive agents derive any benefits from a common pool of funds. However, relying solely on altruistic behavior from constructive agents is insufficient to support the establishment of cooperation in one-shot games. Importantly, the altruistic behavior of constructive agents provides positive incentives for others in society, while the behavior of destructive agents generates negative incentives. By simultaneously considering these positive and negative incentive behaviors balliet2011reward; szolnoki2010reward; szolnoki2017second, the dynamics of cooperation maintenance are altered. Therefore, our study highlights the importance of balancing the interaction of positive and negative incentive behaviors.
@@
In our study, constructive and destructive agents represent two typical orientations towards the interests of others: benefiting others and causing harm to others. However, individual self-interest considerations have been overlooked. Considering that self-interest is a characteristic of rational human behavior, an integration of individual actions build upon two dimensions of behavioral orientation within the social value orientation framework bogaert2008social–preference for one’s own gains and preference for opponent’s gains–will contribute to a more comprehensive understanding of cooperation emergence and promotion under diverse behavioral frameworks. On the other hand, in this challenging one-shot game scenario, cooperation cannot be sustained by relying solely on the constructive agents. In conjunction with reciprocity mechanisms, can the altruistic behaviors exhibited by constructive agents, promote the development of cooperation? Furthermore, by integrating behavioral information transmission, analyzing the biased strategies of constructive and destructive agents towards defectors or cooperators becomes feasible, aiding in a deeper understanding of how diverse behaviors affect the cooperation dynamics. Future research focusing on direct reciprocity, indirect reciprocity, and network reciprocity mechanisms will offer valuable insights into the complex interplay between diverse individual value orientations and the maintenance of cooperation.
Article information
Acknowledgement
@@
We acknowledge the support provided by (i) the National Natural Science Foundation of China (Grant No.11931015, 12271471, 12161089), Major project of National Philosophy and Social Science Foundation of China (Grants No. 22&ZD158 and 22VRCO49) to L.S.; (ii) a JSPS Postdoctoral Fellowship Program for Foreign Researchers (Grant No. P21374), and an accompanying Grant-in-Aid for Scientific Research from JSPS KAKENHI (Grant No. JP 22KF0303) to C.S.; (iii) the grant-in-Aid for Scientific Research from JSPS, Japan, KAKENHI (Grant No. JP 20H02314 and JP 23H03499) awarded to J.T; (iv) China Scholarship Council (Grant No. 202308530309) and Yunnan Provincial Department of Education Science Research Fund Project (Grant No. 2024Y503) to Z.H., and (v) Yunnan Provincial Department of Education Science Research Fund Project (Grant No. 2024Y502) to Y.D.
@@
Conflict of interest
We declare that no conflict of interests.
Appendix A. Finite populations: invasion analysis among cooperation, defection and constructive agents
@@
$\begin{gathered}P_{X}=\pi_{X}(1X,\ (N-1)Y).\end{gathered}$ (A.1)
@@
$\begin{gathered}P_{Y}=\pi_{Y}(1X,(N-1)Y)\cdot p_{1}\\ +\pi_{Y}(0X,NY)\cdot p_{2},\end{gathered}$ (A.2)
@@
$\begin{split}p_{1}&=\frac{\binom{M-2}{N-2}\binom{1}{1}}{\binom{M-1}{N-1}}\\ &=\frac{N-1}{M-1},\end{split}$ (A.3)
$p_{2}$ $=\frac{\binom{M-2}{N-1}\binom{1}{0}}{\binom{M-1}{N-1}}$ (A.4)
@@
$r<N\cdot\frac{M-1}{M-N}.$ (A.6)
In the limit $M\to\infty$ , the above condition is reduced to $r<N$, thereby encompassing the condition for the social dilemma: $0<r<N$. In the subsequent analysis, we will proceed under the assumption that inequality (A.6) holds.
@@
$0<r<1-\frac{d_{2}}{M-1},$ (A.9)
strategy $CA$ can invade the population of strategy $C$ iff condition (A.9) holds.
@@
$r>1-\frac{d_{2}}{M-1},$ (A.10)
when $r$ satisfies condition (A.10), the strategy $CA$ can never invade the population of strategy $C$
@@
$r>1-(N-1)\cdot d_{2},$ (A.12)
strategy $C$ will invade the population of strategy $CA$ iff condition (A.12) holds.
@@
Appendix B. Finite populations: invasion analysis of a simplified three-strategy model involving cooperation, defection and integrated exiters
@@
If $P_{D}>P_{C}$, strategy $D$ successfully invades the population of strategy $C$ , consistent with (A.6), i.e., when
$r<N\cdot\frac{M-1}{M-N}.$ (B.2)
@@
Since inequality (B.2) includes the condition $0<r<N$, it follows that $P_{C}<P_{D}$, indicating that strategy $C$ will never invade the population of strategy $D$.
@@
In the above equation (B.10), if $d_{2}-d_{1}>0$, then $P_{D}>P_{E}$, and strategy $D$ will invade the population of strategy $E$. Conversely, if $d_{2}-d_{1}<0$, then $P_{D}<P_{E}$, and strategy $D$ does not invade the population of strategy $E$. When $d_{2}=d_{1}$, strategy $D$ and strategy $E$ cannot invade each other.
@@
Similarly, in equation (B.11), if $d_{2}-d_{1}<0$, then $P_{E}>P_{D}$, and strategy $E$ will invade the population of strategy $D$. Conversely, if $d_{2}-d_{1}>0$, then $P_{E}<P_{D}$, and strategy $E$ does not invade the population of strategy $D$. When $d_{2}=d_{1}$, strategy $D$ and strategy $E$ cannot invade each other.
@@
However, when $d_{2}-d_{1}=0$, it follows from (B.10) and (B.11) that $PE=PD=0$, indicating that strategies $D$ and $E$ cannot invade each other, and there is no competition between them. For strategies C and E, according to (B.4) and (B.7), we have $P_{C}=r-1$ and $P_{E}=0$. Therefore, when $r>1$, strategy $E$ does not invade the population of strategy $C$; instead, strategy $C$ invades the population of strategy $E$. Additionally, based on (B.1) and (B.3), strategy $D$ always invades the population of strategy $C$. Thus, this leads to $E\rightarrow C\rightarrow D$, resulting in the dominance of strategy $D$. Conversely, when $r<1$, strategy $C$ does not invade strategy $E$, while strategy $E$ invades strategy $C$. Combined with (B.1) and (B.3), since strategy $D$ always invades strategy $C$, both strategies $E$ and $D$ invade strategy $C$, ultimately leading to the coexistence of strategies $E$ and $D$.